### v1

@@ A.3. Stability of the spectral flow
Let $A_{t}$ be a continuous operator family and $E\in\mathbb{R}\setminus\sigma_{\rm ess}(\{A_{t}\})$. There is $\varepsilon>0$ so that, for any norm-resolvent continuous family $(B_{t})$ of self-adjoint operators satisfying $d(A_{t},B_{t})<\varepsilon$ for all $t\in[0,1]$ and $B_{0}=A_{0}$, $B_{1}=A_{1}$, it holds ${\mathrm{Sf}}(A_{t},E,[0,1])={\mathrm{Sf}}(B_{t},E,[0,1])$.
We emphasize that we require that the paths $(A_{t})$ and $(B_{t})$ have the same endpoints here. This is due to the boundary terms in (27).
Proof.
Let $\{t_{i}\}$ be a subdivision for $\{A_{t}\}$, and $a_{i}\geq 0$ the corresponding widths, so that (26) holds. Let us fix $1\leq i\leq M$. For all $t\in[t_{i-1},t_{i}]$, we apply Lemma A.3, and find $\varepsilon(t)>0$ so that the spectral projector on $(E-a_{i},E+a_{i})$ is continuous on $\{B,\ d(A_{t},B)<\varepsilon(t)\}$. We set
$U(t):=\left\{t^{\prime}\in[t_{i-1},t_{i}],\ d(A_{t},A_{t^{\prime}})<\frac{1}{2}\varepsilon(t)\right\}.$
The open sets $U(t)$ cover the compact interval $[t_{i-1},t_{i}]$, so we can find a sub-division $t_{i-1}=t_{i}^{(1)}<t_{i}^{(2)}<\cdots<t_{i}^{(N_{i})}=t_{i}$ so that the union of $\left(U(t_{i}^{(j)})\right)_{1\leq j\leq N_{i}}$ covers $[t_{i-1},t_{i}]$.
We set $\varepsilon_{i}:=\frac{1}{2}\min_{j}\{\varepsilon(t_{i}^{(j)})\}$. Now, assume that $(B_{t})$ is such that $d(A_{t},B_{t})<\varepsilon_{i}$ for all $t\in[t_{i-1},t_{i}]$. Then, for such a $t$, let $1\leq j\leq N_{i}$ be such that $t\in U(t_{i}^{(j)})$, and we have
$d(A_{t_{i}^{(j)}},B_{t})<d(A_{t_{i}^{(j)}},A_{t})+d(A_{t},B_{t})\leq\frac{1}{2}\varepsilon(t_{i}^{(j)})+\varepsilon_{i}<\varepsilon(t_{i}^{(j)}).$
This shows that $P_{(E-a_{i},E+a_{i})}(B_{t})$ is well-defined, with
$\mathrm{rank}\,P_{(E-a_{i},E+a_{i})}(B_{t})=\mathrm{rank}\,P_{(E-a_{i},E+a_{i})}(A_{t}).$
In order to gather all the intervals $[t_{i-1},t_{i}]$, we set $\varepsilon:=\min_{i}\varepsilon_{i}$. For a family $(B_{t})$ with $d(A_{t},B_{t})<\varepsilon$, we deduce that the subdivision $(t_{i})$ with widths $(a_{i})$ is a valid one for $(B_{t})$. In particular, we find that $E+a_{i}$ and $E+a_{i+1}$ are in the resolvent set for all choices of $B_{t_{i}}$ with $d(A_{t_{i}},B_{t_{i}})<\varepsilon$, and hence
$\mathrm{rank}\,P_{(E+a_{i},E+a_{i+1})}(B_{t_{i}})=\mathrm{rank}\,P_{(E+a_{i},E+a_{i+1})}(A_{t_{i}}).$
Since we also have $B_{0}=A_{0}$ and $B_{1}=A_{1}$, we obtain ${\mathrm{Sf}}(A_{t},E,[0,1])={\mathrm{Sf}}(B_{t},E,[0,1])$. ∎
@@
Recall the Weyl’s theorem, which states that if $A$ is self-adjoint and $K$ is compact symmetric, then $A+sK$ is self-adjoint, and $\sigma_{\rm{ess}}(A+sK)=\sigma(A)$ for all $s$: the essential gaps are stable under compact perturbations, and therefore never close. Applying the previous lemma to $A_{t}^{(s)}:=A_{t}+sK$ gives the following result.
@@
${\mathrm{Sf}}(A_{t},E,[0,1])={\mathrm{Sf}}\left((C-A_{t})^{-1},(C-E)^{-1},[0,1]\right)={\mathrm{Sf}}\left((C-A_{t})^{-1}-(C-E)^{-1},0,[0,1]\right)$
@@
A.4. The spectrum quasi–periodic case

### v3

@@ A.3. Stability of the spectral flow
Let $t\mapsto A_{t}$ be a norm-resolvent continuous family of self-adjoint operators, and $E\in\mathbb{R}\setminus\sigma_{\rm ess}(\{A_{t}\})$. There is $\varepsilon>0$ so that, for any norm-resolvent continuous family $(B_{t})$ of self-adjoint operators satisfying $d(A_{t},B_{t})<\varepsilon$ for all $t\in[0,1]$ and $B_{0}=A_{0}$, $B_{1}=A_{1}$, it holds ${\mathrm{Sf}}(A_{t},E,[0,1])={\mathrm{Sf}}(B_{t},E,[0,1])$.
We emphasize that we require the paths $(A_{t})$ and $(B_{t})$ to have the same endpoints here. This is due to the boundary terms in (30). The proof follows the exact same lines as in [31, Proposition 3], so we do not repeat it here.
@@
Recall Weyl’s theorem, which states that if $A$ is self-adjoint and $K$ is compact symmetric, then $A+sK$ is self-adjoint, and $\sigma_{\rm{ess}}(A+sK)=\sigma_{\rm{ess}}(A)$ for all $s$: the essential gaps are stable under compact perturbations, and therefore never close. Applying the previous lemma to $A_{t}^{(s)}:=A_{t}+sK$ gives the following result.
@@
${\mathrm{Sf}}(A_{t},E,[0,1])={\mathrm{Sf}}\left((C-A_{t})^{-1},(C-E)^{-1},[0,1]\right)={\mathrm{Sf}}\left((C-A_{t})^{-1}-(C-E)^{-1},0,[0,1]\right).$
@@
A.4. The translation equivariant case
@@
Acknowledgements
The research leading to these results has received funding from CNRS, AAP IRL 2022, and from ANID, Chile, through the CMM Basal grant FB210005 and through Fondecyt Project 11220194. We thank the anonymous referees for the thorough reading of the manuscript and their suggestions that greatly improved the paper.
Statements
All authors declare that they have no conflicts of interest. The Python code used to generate the figures is available at https://gitlab.com/davidgontier/softwall_jacobimatrix.